OpenStudy (anonymous):

I need help with 3 homework questions. 1. Write 1/3 • 1/4 in lowest terms 2. Write 4/5 • 3/8 in lowest terms 3. Write - 6/11 • 3/4 in lowest terms

OpenStudy (anonymous):

Please Please Please Please PLEASE help me!!!!!!!!!!!!!!!!!

jigglypuff314 (jigglypuff314):

3 * 4 = 12 right? therefore 1/3 * 1/4 = 1/12

jigglypuff314 (jigglypuff314):

for 4/5 • 3/8 4 * 3 = 12 and 5 * 8 = 40 therefore 4/5 • 3/8 = 12/40 but you can figure out that both 12 and 40 can be divided by 4 so 12/4 = 3 and 40/4 = 10 therefore 4/5 • 3/8 = 12/40 = 3/10

OpenStudy (anonymous):

Oh thank you so so so SO much! You're a lifesaver! c:

jigglypuff314 (jigglypuff314):

you're welcome :) do you understand the concept now, or would you still need help on the third one?

OpenStudy (anonymous):

No I understand the concept now. Again thank you!